40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4370     Entry time: Wed Mar 2 22:04:22 2011
Author: Kevin 
Type: Update 
Category: Electronics 
Subject: POY Shot Noise Measurement 

The previous measurement for the shot noise of POY had the dark noise at ~100 nV/rtHz. I redid the measurement and got 26 nV/rtHz for the dark noise. I think that when I made the previous measurement, the spectrum analyzer had automatically added some attenuation to the input that I failed to remove. This added attenuation raised the noise floor of the measurement making the dark noise of POY appear larger than it is.

The updated measurement can be found on the wiki at http://lhocds.ligo-wa.caltech.edu:8000/40m/Electronics/POY.

ELOG V3.1.3-